# Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper

### 11 Mar 2026 v2

@@ Abstract
We present improved lower bounds for five classical Ramsey numbers: $\mathbf{R}(3,13)$ is increased from $60$ to $61$, $\mathbf{R}(3,18)$ from $99$ to $100$, $\mathbf{R}(4,13)$ from $138$ to $139$, $\mathbf{R}(4,14)$ from $147$ to $148$, and $\mathbf{R}(4,15)$ from $158$ to $159$. These results were achieved using AlphaEvolve, an LLM-based code mutation agent. Beyond these new results, we successfully recovered lower bounds for all Ramsey numbers known to be exact, and matched the best known lower bounds across many other cases. These include bounds for which previous work does not detail the algorithms used.
@@ 1.1 Summary of Our Results
In this paper, we focus on small Ramsey numbers [13] where $r,s\leq 22$. Our main results are that we improve $\mathbf{R}(3,13)$ from $60$ [10] to $61$, $\mathbf{R}(3,18)$ from $99$ [4] to $100$, $\mathbf{R}(4,13)$ from $138$ [3] to $139$, $\mathbf{R}(4,14)$ from $147$ [3] to $148$, and $\mathbf{R}(4,15)$ from $158$ [15] to $159$. Furthermore, we match the SoTA on many other cells in Table 1, including all cells for which the exact value of $\mathbf{R}(r,s)$ is known. The search algorithms used to match or improve these 28 bounds can be categorized into four distinct families of initialization; these are summarized in Table 2 and detailed in the Appendix. Note that for some of these cells, prior work only exhibits the graph implying the bound, without giving the search algorithm that discovered the graph [13].
Note: Our constructions for the the cells $\mathbf{R}(3,13),\mathbf{R}(3,18),\mathbf{R}(4,13),\mathbf{R}(4,14)$ and $\mathbf{R}(4,15)$ are made public on Github at this link.
@@ 2 AlphaEvolve Meta-search, and Discovered Search Algorithms
AlphaEvolve [12, 14] maintains a family of search algorithms (that search for graphs that establish Ramsey lower bounds), using an LLM to evolve them to find better search algorithms. Typically this process uses a synthetic objective function to guide AlphaEvolve, as a proxy for the true objective (the size of the graph being discovered). This detail will become evident in our presentation of Algorithm 1 below.
@@ 16 $\mathcal{P}\leftarrow\mathcal{P}\cup\{(\texttt{p}_{\texttt{new}},S)\}$
Algorithms Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working PaperReinforced Generation of Combinatorial Structures: Ramsey Numbers Working PaperReinforced Generation of Combinatorial Structures: Ramsey Numbers Working PaperReinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper and Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper detail the five search algorithms that AlphaEvolve found for the cells $\mathbf{R}(3,13)$, $\mathbf{R}(3,18)$, $\mathbf{R}(4,13)$, $\mathbf{R}(4,14)$, and $\mathbf{R}(4,15)$ respectively. The detailed algorithms for all other colored cells in Table 1 are presented in Algorithms Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working PaperReinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper. Note that each of the search algorithms seems specific to the corresponding cells; their success does not seem to transfer between cells, even when run for a reasonably long time. In Table 2, we summarize the initialization strategies used in the search algorithms. We left out algorithms for $\mathbf{R}(3,3)$ and $\mathbf{R}(3,4)$ because AlphaEvolve just quoted the corresponding adjacency matrices from the LLM’s memory. The search algorithms for each of the cells (while having some differences) are different variants of stochastic search.
@@ 3 Comparison to Prior Work
  • $\mathbf{R}(4,13),\mathbf{R}(4,14),\mathbf{R}(4,15)$, and $\mathbf{R}(6,7)$: Except for $\mathbf{R}(4,15)$, the previous SoTA for these cells were established in [3]. For $\mathbf{R}(4,15)$, the prior SoTA was established in [15]. We improved the bound for $\mathbf{R}(4,13)$ (and $\mathbf{R}(4,14)$), while matching the bound for $\mathbf{R}(6,7)$. Notably, the search algorithms discovered by AlphaEvolve for $\mathbf{R}(4,13)$ and $\mathbf{R}(6,7)$ (Algorithms Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper and Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper) initialize with the same algebraic structures—cubic residue and Paley graphs, respectively—as those used in [3]. For $\mathbf{R}(4,14)$, Exoo [3] initialized the search with a cubic graph, whereas AlphaEvolve (Algorithm Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper) used a class of circulant graphs with varying periodicities. The subsequent optimization strategies for all the cells differ significantly: AlphaEvolve utilizes evolved heuristics rather than the standard simulated annealing employed in [3].
@@
    Our approach for $\mathbf{R}(4,15)$ (Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper) diverges from the Cayley graph initialization of [3] (which is the SoTA prior to [15]), starting instead from a Paley graph. We unfortunately could not compare to [15] as it is referred to as personal communication in [13].
@@
  • $\mathbf{R}(4,6)$: We match the SoTA for this cell, which was achieved by [5]. While [5] initializes the search with a random graph, AlphaEvolve opts for a cyclic graph initialization. Despite this difference, the subsequent search procedure discovered by AlphaEvolve (Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper) bears structural similarities to the heuristic search described in [5].
@@
  • $\mathbf{R}(4,10)$: The current SoTA for this cell is attributed to [9], who employed a branch-and-bound search on circulant colorings. In contrast, the algorithm discovered by AlphaEvolve (Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper) initiates by sampling from a distribution over circulant graphs, followed by a custom local search that integrates sophisticated tabu search mechanisms with sequential growth.
@@
This comparison demonstrates that for many cells — even those where AlphaEvolve only matched the SoTA — AlphaEvolve’s search algorithms differ materially from prior methods, despite often sharing initialization families (e.g., algebraic or cyclic graphs). At a meta-level, the algorithms generated by AlphaEvolve exhibit three salient properties: a) they exploit known algebraic structures relevant to the specific cell; b) unlike many human-designed algorithms, they typically chain multiple heuristics or execute them in parallel; and c) they incorporate custom heuristics for approximate clique and independent set counting to accelerate candidate evaluation. Furthermore, for cases such as $\mathbf{R}(4,10)$ (see Reinforced Generation of Combinatorial Structures: Ramsey Numbers Working Paper), AlphaEvolve’s search strategies appear novel and, to the best of our knowledge, are absent from existing literature.

### 21 Apr 2026 v5

@@ Abstract
We present improved lower bounds for nine classical Ramsey numbers: $\mathbf{R}(3,13)$ is increased from $60$ to $61$, $\mathbf{R}(3,18)$ from $99$ to $100$, $\mathbf{R}(4,13)$ from $138$ to $139$, $\mathbf{R}(4,14)$ from $147$ to $148$, $\mathbf{R}(4,15)$ from $158$ to $159$, $\mathbf{R}(4,16)$ from $170$ to $174$, $\mathbf{R}(4,18)$ from $205$ to $209$, $\mathbf{R}(4,19)$ from $213$ to $219$, and $\mathbf{R}(4,20)$ from $234$ to $237$. These results were achieved using AlphaEvolve, an LLM-based code mutation agent. Beyond these new results, we successfully recovered lower bounds for all Ramsey numbers known to be exact, and matched the best known lower bounds across many other cases. These include bounds for which previous work does not detail the algorithms used.
@@ 1.1 Summary of Our Results
In this paper, we focus on small Ramsey numbers [13] where $r,s\leq 22$. Our main results are that we improve $\mathbf{R}(3,13)$ from $60$ [10] to $61$, $\mathbf{R}(3,18)$ from $99$ [4] to $100$, $\mathbf{R}(4,13)$ from $138$ [3] to $139$, $\mathbf{R}(4,14)$ from $147$ [3] to $148$, $\mathbf{R}(4,15)$ from $158$ [15], $\mathbf{R}(4,16)$ from $170$ [15] to $174$, $\mathbf{R}(4,18)$ from $205$ [13] to $209$, $\mathbf{R}(4,19)$ from $213$ [13] to $219$, and $\mathbf{R}(4,20)$ from $234$ [13] to $237$. Furthermore, we match the SoTA on many other cells in Table 1, including all cells for which the exact value of $\mathbf{R}(r,s)$ is known. The search algorithms used to match or improve these 28 bounds can be categorized into four distinct families of initialization; these are summarized in Table 2 and detailed in the Appendix. Note that for some of these cells, prior work only exhibits the graph implying the bound, without giving the search algorithm that discovered the graph [13].
Note: Our constructions for the the cells $\mathbf{R}(3,13),\mathbf{R}(3,18),\mathbf{R}(4,13),\mathbf{R}(4,14)$, $\mathbf{R}(4,15)$, $\mathbf{R}(4,16)$, $\mathbf{R}(4,18)$, $\mathbf{R}(4,19)$, and $\mathbf{R}(4,20)$ are made public on Github at this link. Additionally, the search algorithms discovered by AI are also made public on Github at this link.
@@ 2 AlphaEvolve Meta-search, and Discovered Search Algorithms
AlphaEvolve [12, 14] can be used to maintain a family of search algorithms (that search for graphs that establish Ramsey lower bounds), using an LLM to evolve them to find better search algorithms. Typically this process uses a synthetic objective function to guide AlphaEvolve, as a proxy for the true objective (the size of the graph being discovered). This detail will become evident in our presentation of Algorithm 1 below.
@@ 16 $\mathcal{P}\leftarrow\mathcal{P}\cup\{(\texttt{p}_{\texttt{new}},S)\}$
Algorithms Reinforced Generation of Combinatorial Structures: Ramsey NumbersReinforced Generation of Combinatorial Structures: Ramsey NumbersReinforced Generation of Combinatorial Structures: Ramsey NumbersReinforced Generation of Combinatorial Structures: Ramsey NumbersReinforced Generation of Combinatorial Structures: Ramsey NumbersReinforced Generation of Combinatorial Structures: Ramsey NumbersReinforced Generation of Combinatorial Structures: Ramsey NumbersReinforced Generation of Combinatorial Structures: Ramsey Numbers, and Reinforced Generation of Combinatorial Structures: Ramsey Numbers detail the nine search algorithms that AlphaEvolve found for the cells $\mathbf{R}(3,13)$, $\mathbf{R}(3,18)$, $\mathbf{R}(4,13)$, $\mathbf{R}(4,14)$, and $\mathbf{R}(4,15)$, $\mathbf{R}(4,16)$, $\mathbf{R}(4,18)$, $\mathbf{R}(4,19)$ and $\mathbf{R}(4,20)$ respectively. The detailed algorithms for all other colored cells in Table 1 are presented in Algorithms Reinforced Generation of Combinatorial Structures: Ramsey NumbersReinforced Generation of Combinatorial Structures: Ramsey Numbers. Note that each of the search algorithms seems specific to the corresponding cells; their success does not seem to transfer between cells, even when run for a reasonably long time. In Table 2, we summarize the initialization strategies used in the search algorithms. We left out algorithms for $\mathbf{R}(3,3)$ and $\mathbf{R}(3,4)$ because AlphaEvolve just quoted the corresponding adjacency matrices from the LLM’s memory. The search algorithms for each of the cells (while having some differences) are different variants of stochastic search.
@@ 3 Comparison to Prior Work
  • $\mathbf{R}(4,13),\mathbf{R}(4,14),\mathbf{R}(4,15)$$\mathbf{R}(4,15)$, $\mathbf{R}(4,16)$, $\mathbf{R}(4,18)$ and $\mathbf{R}(6,7)$: Except for $\mathbf{R}(4,s),s\in\{15,16,18\}$, the previous SoTA for these cells were established in [3]. For $\mathbf{R}(4,15)$ and $\mathbf{R}(4,16)$ the prior SoTA was established in [15]. We improved the bound for $\mathbf{R}(4,13)$ (and $\mathbf{R}(4,14)$), while matching the bound for $\mathbf{R}(6,7)$. Notably, the search algorithms discovered by AlphaEvolve for $\mathbf{R}(4,13)$ and $\mathbf{R}(6,7)$ (Algorithms Reinforced Generation of Combinatorial Structures: Ramsey Numbers and Reinforced Generation of Combinatorial Structures: Ramsey Numbers) initialize with the same algebraic structures—cubic residue and Paley graphs, respectively—as those used in [3]. For $\mathbf{R}(4,14)$, Exoo [3] initialized the search with a cubic graph, whereas AlphaEvolve (Algorithm Reinforced Generation of Combinatorial Structures: Ramsey Numbers) used a class of circulant graphs with varying periodicities. The subsequent optimization strategies for all the cells differ significantly: AlphaEvolve utilizes evolved heuristics rather than the standard simulated annealing employed in [3]. For $\mathbf{R}(4,18)$, the previous SoTA [13] was obtained via differencing between other Ramsey cells. However, AlphaEvolve developed a new ground-up heuristic for this cell.
@@
    Our approach for $\mathbf{R}(4,15)$ (Reinforced Generation of Combinatorial Structures: Ramsey Numbers) diverges from the Cayley graph initialization of [3] (which is the SoTA prior to [15]), starting instead from a Paley graph. We unfortunately could not compare to [15] as it is referred to as personal communication in [13].
@@
  • $\mathbf{R}(4,6)$: We match the SoTA for this cell, which was achieved by [5]. While [5] initializes the search with a random graph, AlphaEvolve opts for a cyclic graph initialization. Despite this difference, the subsequent search procedure discovered by AlphaEvolve (Reinforced Generation of Combinatorial Structures: Ramsey Numbers) bears structural similarities to the heuristic search described in [5].
@@
  • $\mathbf{R}(4,10)$: The current SoTA for this cell is attributed to [9], who employed a branch-and-bound search on circulant colorings. In contrast, the algorithm discovered by AlphaEvolve (Reinforced Generation of Combinatorial Structures: Ramsey Numbers) initiates by sampling from a distribution over circulant graphs, followed by a custom local search that integrates sophisticated tabu search mechanisms with sequential growth.
@@
This comparison demonstrates that for many cells — even those where AlphaEvolve only matched the SoTA — AlphaEvolve’s search algorithms differ materially from prior methods, despite often sharing initialization families (e.g., algebraic or cyclic graphs). At a meta-level, the algorithms generated by AlphaEvolve exhibit three salient properties: a) they exploit known algebraic structures relevant to the specific cell; b) unlike many human-designed algorithms, they typically chain multiple heuristics or execute them in parallel; and c) they incorporate custom heuristics for approximate clique and independent set counting to accelerate candidate evaluation. Furthermore, for cases such as $\mathbf{R}(4,10)$ (see Reinforced Generation of Combinatorial Structures: Ramsey Numbers), AlphaEvolve’s search strategies appear novel and, to the best of our knowledge, are absent from existing literature.